OpenStudy (anonymous):

How would you solve q + 1/3 > 1/2?

OpenStudy (anonymous):

\[q+\frac{ 1 }{3 }>\frac{ 1 }{ 2 }\]\[q>\frac{ 1 }{ 2 }-\frac{ 1 }{ 3 }\]\[q>\frac{ 3 }{ 6 }-\frac{ 2 }{ 6 }\]\[q>\frac{ 1 }{ 6 }\]